Using identities, evaluate. 
99^2 ?????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

99^2
= (100 -1)2
=100^2-200 +1^2
=10000-200+1
= 9801

Thanks